# EvTTC: An Event Camera Dataset for Time-to-Collision Estimation

### v1

@@ I Introduction
FCWD (Ours) 1920$\times$1200 [160-295] 1280$\times$720 [99-197] TTC, Depth, RTK GPS
@@
Event-based cameras, inspired by the biological mechanisms of the human visual system, operate differently from traditional frame-based cameras. Rather than capturing entire frames, event-based cameras asynchronously report changes in brightness at individual pixels. This approach offers spatio-temporal sparsity, microsecond-level temporal resolution, and high dynamic range, making event-based cameras well-suited for perception tasks involving rapid motion, such as robotics and autonomous driving [21][22][23][24][25].
@@
Outline: The rest of the paper is organized as follows. First, a literature review on datasets of event cameras and those for TTC tasks is provided in Sec. II. Then we provide details of our sensor setup and calibration in Sec. III, followed with descriptions of all the sequences (Sec. IV). Moreover, the design of the small-scale TTC testbed is disclosed in Sec. IV. Finally, we establish the TTC benchmark by providing a comprehensive experiment results in Sec. VI, and make the conclusion in Sec. VII.
@@ II Related Work
Over the past decade, numerous datasets have been developed to address specific tasks in autonomous driving. Our literature review focuses on two categories: 1) datasets for predicting a traffic collision and 2) outdoor datasets collected using event cameras. Table I provides an overview of these datasets, briefly showcasing the different sensor configurations they use and their different focuses.
The first category focuses on collision risk prediction and TTC estimation. The DR(eye)VE Project [1] and DADA-2000 [2] are notable for their rich data and diverse scenarios. DR(eye)VE provides over 500k frames of normal driving scenes along with driver eye-tracking data, and DADA-2000 includes 2k emergency braking sequences and eye-tracking data from volunteers observing accident videos. These datasets facilitate predicting the risk of a collision by using drivers’ attention data as ground truth. Moreover, [3] provides manually labeled 2D bounding boxes and collision labels, featuring 122 emergency braking scenes and 100 normal driving scenes from YouTube. However, as shown in Tab. I, these datasets lack ground truth of depth and trajectory, limiting their usage in developing TTC estimation algorithms. Among these datasets, the TSTTC [4] dataset is closest to our focus, particularly collected for TTC estimation in driving scenarios. TSTTC is a large-scale dataset with over 200k frames captured from real-world highway and urban environments using monocular cameras with three different focal lengths, covering a detection range of 16-96 meters. It includes ground-truth TTC and 2D/3D bounding boxes, which are crucial for TTC estimation. However, TSTTC mainly focuses on normal driving scenarios with low relative speed between vehicles, and lacks emergency braking under high-speed scenarios, which are essential for evaluating TTC algorithms in extreme conditions.
@@
The second category is those collected using event cameras, which leverages the specialty of event cameras to handle high-speed motion and challenging lighting conditions. Among them, MVSEC [5] is the first stereo event camera dataset, where the stereo rig is handheld or mounted on a mobile platform, such as a drone, an automobile and a motorcycle. Afterwards, Gehrig et al.[6] present a large-scale stereo event dataset with higher spatial resolution, better sensor synchronization, and a wide baseline, leading to better depth estimation accuracy. To handle extreme lighting conditions, ViViD++[7] integrates thermal sensors alongside event cameras, but hardware triggered synchronization is not witnessed. Limited by the immaturity of event cameras at that time, high definition (HD) data are not available in these datasets until the appearance of M3ED[8], which utilizes a Prophesee EVKv4 event camera with a spatial resolution of 1280$\times$720 pixels. In general, these datasets lack specific scenarios and data for the development of TTC tasks.
@@ III-A Sensors
8-mm Lens Camera Pair Prophesee EVKv4 FLIR Blackfly S 1280$\times$720 1/2.5” FoV: 41°$\times$24 1920$\times$1200 1/2.3” FoV: 44°$\times$29° Rate: 20Hz
16-mm Lens Camera Pair Prophesee EVKv4 FLIR Blackfly S 1280$\times$720 1/2.5” FoV: 22°$\times$12° 1920$\times$1200 1/2.3” FoV: 23°$\times$14° Rate: 20Hz
@@
Figure 2: Illustration of our synchronization scheme.
@@
Figure 4: The top-view schematic of the dataset scenarios. The arrow represents the direction of movement. The lateral shadowing of the DCV in the CCRs and CCRm scenarios indicates that data are collected across different lane positions.
@@
The raw data used for estimating TTC contain RGB images and corresponding event streams. To cover a relatively complete sensing range, we employ two groups of RGB-Event camera pair equipped with lens of different focal length. The camera pair using 8-mm lens consists of an RGB camera and an event camera, and it covers the close sensing range (see A in Fig. LABEL:fig:eye_catcher). As a complementary, the camera pair using 16-mm lens is used to cover the distant sensing range (see B in Fig. LABEL:fig:eye_catcher). In each pair, the event camera and the RGB camera are rigidly attached with a narrow baseline of 4 cm, and thus, the extrinsic parameters (w.r.t the coordinate system of the sensor suite) can be approximately shared, simplifying data fusion across the two sensor modalities. The employed event camera is Prophesee EVK-v4 (sensor size 1/2.5”) that has a spatial resolution of $1280\times 720$ pixels. To ensure a similar field of view (FOV), an FLIR Blackfly-S RGB camera (sensor size 1/2.3”) with a $1920\times 1200$ resolution is used and equipped with an identical lens. Detailed parameters of each device can be found in Tab. II.
The ground-truth data consists of depth information observed from the host vehicle and the involved two vehicles’ motion status (e.g., position, speed, and orientation, etc). To provide these information, we employ a solid-state LiDAR and two high-performance GNSS/INS units (see Tab. II). A Livox HAP LiDAR is rigidly attached to the two camera pairs and used for collecting ground-truth depth. As for the GNSS/INS devices, a UG005 X1 device is used on the host vehicle, and a CGI-610 device is used on the lead vehicle, both for providing high-frequency motion status of the platform. Note that international organizations of vehicle safety assessment, such as Euro New Car Assessment Program (NCAP), require the motion status of involved vehicles to be collected at a frequency of no lower than 100 Hz. To this end, both GNSS/INS units are configured to provide motion status at 100 Hz.
@@ III-B Time Synchronization
We synchronize all sensors in the sensor suite at the hardware level. To this end, we utilize Precision Timing Protocol (PTP) [26] and the generalized Precision Time Protocol (gPTP) [27], which can provide sub-microsecond synchronization accuracy in the Ethernet.
The time synchronization among all the devices is illustrated in Fig. 2. First, we use a PTP server to synchronize GNSS’s clock with the system clock of the industrial personal computer (IPC). The IPC is then designated as the master clock, using the PTP protocol to synchronize with the RGB cameras. Since the Livox HAP only supports the gPTP protocol, we use gPTP to synchronize between the IPC and the LiDAR. To synchronize the event cameras with the system time, we utilize a micro-controller to generate four synchronization pulses at 20 Hz. These pulses are used to simultaneously trigger the two RGB cameras and the two event cameras. The RGB and event cameras are configured in external trigger mode. When the RGB camera receives a synchronization pulse, it begins to capture an image and record its timestamp. Meanwhile, the event camera would also be triggered by the synchronization pulse, immediately generating a signal with a timestamp. The temporal offset between the RGB camera and the event camera can be easily calculated as the difference between the two timestamps. In addition, the integrated navigation systems UG005 X1 and CGI-610 are synchronized via their built-in GNSS clocks.
@@ III-C Calibration
We discuss all involved calibration tasks in this part, including the cameras’ intrinsic and extrinsic calibration, extrinsic calibration between cameras and the IMU, and that between cameras and the LiDAR. The precise CAD model of our sensor suite is illustrated in Fig. 3. All calibration results are made available in both YAML and HDF5 formats.
@@ IV Sequences Overview
This section presents a comprehensive overview of the dataset. Following the Euro NCAP AEB test protocol, we collect data for Car-to-Car and Car-to-Pedestrian emergency collision scenarios under various motion conditions (Sec. IV-A), as illustrated in Fig. 4 and Fig. 5. The ground-truth TTC, vehicle poses, and depth information are created by the data from LiDAR and GNSS/INS units (Sec. IV-B).
@@ IV-A1 Car-to-Car Rear scenarios
We collect data from two types of scenarios: Car-to-Car Rear stationary (CCRs) and Car-to-Car Rear moving (CCRm). We use two types of real vehicles and a 1:1 scale inflatable vehicle as collision targets, with the latter intended to simulate more realistic collision scenarios while ensuring safety. In the CCRs, when the Global Vehicle Target (GVT) is a real vehicle, it is placed only longitudinally, as shown in Fig. 5LABEL:sub@fig:ccrs-2 and Fig. 5LABEL:sub@fig:ccrs-3. When the GVT is an inflatable vehicle, it is placed both longitudinally and laterally, as illustrated in Fig. 5LABEL:sub@fig:ccrs-1 and Fig. 5LABEL:sub@fig:ccrs-side. In the CCRm, the GVT moves at a constant speed of 20 km/h (see Fig. 5LABEL:sub@fig:ccrs-m). Additionally, in the CCRs-1 (Fig. 5LABEL:sub@fig:ccrs-1) and CCRm, the Data Collection Vehicle (DCV) and the GVT occupy different lane positions, including fully aligned, partially offset, and completely offset cases. In all scenarios, the DCV approaches the GVT at speeds ranging from approximately 20 to 60 km/h, as detailed in Fig. 6 and Tab. III.
@@ IV-A2 Car-to-Pedestrian scenarios
We collect data from three types of scenarios: Car-to-Pedestrian Longitudinal Adult (CPLA), Car-to-Pedestrian Nearside Adult (CPNA), and Car-to-Pedestrian Nearside Adult Obstructed (CPNAO). We select an inflatable dummy as the Adult Pedestrian Target (APT). In the CPLA, the APT is placed in the middle of the road, facing away from the DCV, as shown in Fig. 5LABEL:sub@fig:cpla. In the CPNA, the APT crosses the road at a constant speed, as illustrated in Fig. 5LABEL:sub@fig:cpna. In the CPNAO, the APT suddenly crosses the road from outside the DCV’s field of view (see Fig. 5LABEL:sub@fig:cpnao). In all scenarios, the DCV approaches the APT at speeds ranging from approximately 20 to 60 km/h, as shown in Fig. 6 and Tab. III.
@@ IV-B2 Pose and Depth
We use the LiDAR odometry FAST-LIO2 [30] to obtain ground-truth poses and velocity-compensated point clouds. For each pose, we extract the relative poses of several preceding and succeeding point cloud frames to generate local point clouds. Next, the Hidden Point Removal (HPR) operator [31] is applied to eliminate occlusion effects in the local point clouds. Finally, the local point clouds are projected into the RGB camera coordinate frames of the 8-mm and 16-mm lens camera pairs to obtain ground-truth depth.
@@ V SMALL-SCALE TESTBED
Hybird Optical System Inivation DVXplorer DAHENG MER2 640$\times$480 1/3.5” FoV: 20°$\times$15° 1440$\times$1080 FoV: 17°$\times$13° 1/2.9” Rate: 25Hz
@@
Method CCRs1-low CCRs1-medium CCRs1-high CCRs2-low CCRs2-medium $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) STRTTC[24] 7.54 0.031 9.64 0.026 7.78 0.045 8.15 0.024 11.83 0.024 CMax[32] 2.56 3.833 3.44 3.701 7.39 3.228 5.04 3.246 11.20 3.449 ETTCM[33] 45.11 0.081 44.94 0.081 43.03 0.300 48.83 0.350 52.84 0.098 FAITH[34] 15.15 0.156 20.63 0.155 19.51 0.337 25.83 0.169 47.45 0.191 AEB-Tracker[35] 39.57 $\mathbf{4.2\times 10^{-5}}$ 42.41 $\mathbf{3.9\times 10^{-5}}$ 39.04 $\mathbf{1.1\times 10^{-5}}$ 37.09 $\mathbf{9.4\times 10^{-6}}$ 43.17 $\mathbf{7.3\times 10^{-6}}$ Image’s FoE[20] 5.37 0.036 5.12 0.031 1.86 0.041 3.76 0.031 3.85 0.031 Method CCRs2-high CCRm-low CCRm-medium Slider-750 Slider-1000 $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) STRTTC[24] 13.68 0.028 9.87 0.031 10.36 0.023 8.95 0.015 12.43 0.016 CMax[32] 11.10 3.826 9.27 3.658 14.73 3.657 4.16 0.85 2.74 0.93 ETTCM[33] 49.46 0.127 71.05 0.059 58.00 0.121 18.99 0.498 15.20 0.191 FAITH[34] 49.62 0.165 75.78 0.116 56.26 0.11 34.45 0.186 47.94 0.233 AEB-Tracker[35] 42.43 $\mathbf{6\times 10^{-6}}$ 26.95 $\mathbf{1\times 10^{-5}}$ 35.99 $\mathbf{1.1\times 10^{-5}}$ 157.88 $\mathbf{1\times 10^{-5}}$ 165.56 $\mathbf{7.8\times 10^{-6}}$ Image’s FoE[20] 2.92 0.031 5.60 0.027 3.86 0.026 11.35 0.013 7.55 0.012

### v3

@@ I Introduction
EvTTC (Ours) 1920$\times$1200 [160-295] 1280$\times$720 [99-197] TTC, Depth, RTK GPS
@@
Event-based cameras, inspired by the biological mechanisms of the human visual system, operate differently from traditional frame-based cameras. Rather than capturing entire frames, event-based cameras asynchronously report changes in brightness at individual pixels. This approach offers spatio-temporal sparsity, microsecond-level temporal resolution, and high dynamic range, making event-based cameras well-suited for perception tasks involving rapid motion, such as robotics and autonomous driving [21][22][23][24].
@@ II Related Work
In this section, our literature review focuses on three aspects: 1) event-based TTC estimation methods, 2) autonomous driving datasets focused on collision risk prediction and TTC estimation tasks, and 3) event camera-centric datasets. Table I provides an overview of these datasets, briefly showcasing the different sensor configurations they use and their different focuses.
1) Event-based TTC Estimation Methods: Inspired by methods using standard cameras [18][20], most event-based approaches [25][26][27] are built on optical flow derived from event data. Among these methods, [25] and [27] typically use the estimated optical flow to select the Focus of Expansion (FoE) located in the negative half-plane for TTC estimation. However, the performance of these methods is significantly influenced by the accuracy of the estimated FoE. Gallego et al[28] propose the CMax, which can implicitly handle the data association between events. This approach enables precise estimation of TTC through geometric model fitting (e.g., an affine model). To address the event collapse issue in CMax, Shiba et al[29] present a regularization method based on the CMax framework. ETTCM [30] extends the DMin framework [31] to jointly estimate the scene’s inverse depth and global motion to generate the TTC map. Further advancing the field, Li et al[32] propose a two-stage method that first employs a robust linear solver based on a time-variant affine model, followed by further refinement of the resulting model through spatio-temporal registration. In contrast to the aforementioned methods, AEB-Tracker [33] uses an extended Kalman filter to track the state of event blobs (e.g., the tail lights of vehicles), thereby obtaining the scale variation of the target object.
2) Autonomous Driving Datasets Focused on Collision Risk Prediction and TTC Estimation Tasks: Currently, numerous datasets exist for collision risk prediction, such as DR(eye)VE [1] and DADA-2000 [2], both of which offer data from a variety of scenarios. For example, DR(eye)VE provides over 500k frames of normal driving scenes with driver eye-tracking data. Moreover, [3] offers 2D bounding boxes and collision labels for emergency braking and normal driving scenes. However, these datasets lack ground truth of depth and trajectory, limiting their usage for TTC estimation. The TSTTC [4] dataset, which focuses on TTC estimation, contains over 200k frames with ground-truth TTC and bounding boxes. Nevertheless, these datasets only include scenarios under normal driving conditions and lack high-speed emergency braking situations, which are crucial for testing TTC algorithms under extreme conditions.
@@
3) Event Camera-Centric Datasets: There are a number of event camera-centric datasets that include driving scenarios. Among them, MVSEC [5] is the first stereo event camera dataset, where the stereo rig is handheld or mounted on a mobile platform, such as a drone, an automobile and a motorcycle. Afterwards, Gehrig et al.[6] present a large-scale stereo event dataset with higher spatial resolution, better sensor synchronization, and a wide baseline, leading to better depth estimation accuracy. To handle extreme lighting conditions, ViViD++[7] integrates thermal sensors alongside event cameras, but hardware triggered synchronization is not witnessed. Limited by the immaturity of event cameras at that time, high definition (HD) data are not available in these datasets until the appearance of M3ED[8], which utilizes a Prophesee EVK4 event camera with a spatial resolution of 1280$\times$720 pixels. In general, these datasets lack specific scenarios and data for the development of TTC tasks.
@@ III-A Sensors
8-mm Lens Camera Pair Prophesee EVK4 FLIR Blackfly S 1280$\times$720 1/2.5” FoV: 41°$\times$24 1920$\times$1200 1/2.3” FoV: 44°$\times$29° Rate: 20Hz
16-mm Lens Camera Pair Prophesee EVK4 FLIR Blackfly S 1280$\times$720 1/2.5” FoV: 22°$\times$12° 1920$\times$1200 1/2.3” FoV: 23°$\times$14° Rate: 20Hz
@@
Figure -1: Illustration of our synchronization scheme.
@@
Figure 1: The top-view schematic of the dataset scenarios. The arrow represents the direction of movement. The lateral shadowing of the DCV in the CCRs and CCRm scenarios indicates that data are collected across different lane positions. The objects in the scene include the Global Vehicle Target (GVT), Adult Pedestrian Target (APT), and Data Collection Vehicle (DCV).
@@
The raw data used for estimating TTC contain RGB images and corresponding event streams. To cover a relatively complete sensing range, we employ two groups of RGB-Event camera pair equipped with lens of different focal length. The camera pair using 8-mm lens consists of an RGB camera and an event camera, and it covers the close sensing range (see A in Fig. LABEL:fig:eye_catcher). As a complementary, the camera pair using 16-mm lens is used to cover the distant sensing range (see B in Fig. LABEL:fig:eye_catcher). In each pair, the event camera and the RGB camera are rigidly attached with a narrow baseline of 4 cm, and thus, the extrinsic parameters (w.r.t the coordinate system of the sensor suite) can be approximately shared, simplifying data fusion across the two sensor modalities. Detailed parameters of each device can be found in Tab. II.
The ground-truth data consists of depth information observed from the host vehicle and the involved two vehicles’ motion status (e.g., position, speed, and orientation, etc). To provide these information, we employ a solid-state LiDAR and two high-performance GNSS/INS units (see Tab. II). Note that international organizations of vehicle safety assessment, such as Euro New Car Assessment Program (NCAP), require the motion status of involved vehicles to be collected at a frequency of no lower than 100 Hz. To this end, both GNSS/INS units are configured to provide motion status at 100 Hz.
@@ III-B Time Synchronization
To this end, we utilize Precision Timing Protocol (PTP) [34] and the generalized Precision Time Protocol (gPTP) [35], which can provide sub-microsecond synchronization accuracy in the Ethernet.
The time synchronization among all the devices is illustrated in Fig. -1. First, we use a PTP server to synchronize GNSS’s clock with the system clock of the industrial personal computer (IPC). The IPC is then designated as the master clock, using the PTP protocol to synchronize with the RGB cameras. Since the Livox HAP only supports the gPTP protocol, we use gPTP to synchronize between the IPC and the LiDAR. To synchronize the event cameras with the system time, we utilize a micro-controller to generate four synchronization pulses at 20 Hz. These pulses are used to simultaneously trigger the two RGB cameras and the two event cameras. In addition, the integrated navigation systems UG005 X1 and CGI-610 are synchronized via their built-in GNSS clocks.
@@ III-C Calibration
We discuss all involved calibration tasks in this part, including the cameras’ intrinsic and extrinsic calibration, extrinsic calibration between cameras and the IMU, and that between cameras and the LiDAR. The precise CAD model of our sensor suite is illustrated in Fig. III-A. All calibration results are made available in both YAML and HDF5 formats.
@@ IV Sequences Overview
This section presents a comprehensive overview of the dataset. The scenario configurations and vehicle motion conditions are developed based on the Car-to-Car and Car-to-Pedestrian emergency collision scenarios defined in the Euro NCAP AEB test protocol [38]. For further details, please refer to Sec. IV-A and the illustrations in Fig.1 and Fig. 2. The ground-truth TTC, vehicle poses, and depth information are generated by the data from LiDAR and GNSS/INS units (Sec. IV-B).
@@ IV-A1 Car-to-Car Rear scenarios
We collect data from two types of scenarios: Car-to-Car Rear stationary (CCRs) and Car-to-Car Rear moving (CCRm). We use two types of real vehicles and a 1:1 scale inflatable vehicle as collision targets, with the latter intended to simulate more realistic collision scenarios while ensuring safety. In the CCRs, when the GVT is a real vehicle, it is placed only longitudinally, as shown in Fig. 2LABEL:sub@fig:ccrs-2 and Fig. 2LABEL:sub@fig:ccrs-3. When the GVT is an inflatable vehicle, it is placed both longitudinally and laterally, as illustrated in Fig. 2LABEL:sub@fig:ccrs-1 and Fig. 2LABEL:sub@fig:ccrs-side. In the CCRm, the GVT moves at a constant speed of 20 km/h (see Fig. 2LABEL:sub@fig:ccrs-m). Additionally, in the CCRs-1 (Fig. 2LABEL:sub@fig:ccrs-1) and CCRm, the DCV and the GVT occupy different lane positions, including fully aligned, partially offset, and completely offset cases. In all scenarios, the DCV approaches the GVT at speeds ranging from approximately 20 to 60 km/h, as detailed in Fig. 3 and Tab. III.
@@ IV-A2 Car-to-Pedestrian scenarios
We collect data from three types of scenarios: Car-to-Pedestrian Longitudinal Adult (CPLA), Car-to-Pedestrian Nearside Adult (CPNA), and Car-to-Pedestrian Nearside Adult Obstructed (CPNAO). We select an inflatable dummy as the APT. In the CPLA, the APT is placed in the middle of the road, facing away from the DCV, as shown in Fig. 2LABEL:sub@fig:cpla. In the CPNA, the APT crosses the road at a constant speed, as illustrated in Fig. 2LABEL:sub@fig:cpna. In the CPNAO, the APT suddenly crosses the road from outside the DCV’s field of view (see Fig. 2LABEL:sub@fig:cpnao). In all scenarios, the DCV approaches the APT at speeds ranging from approximately 20 to 60 km/h, as shown in Fig. 3 and Tab. III.
@@ IV-B2 Pose and Depth
We use the LiDAR odometry FAST-LIO2 [39] to obtain ground-truth poses and velocity-compensated point clouds. For each pose, we extract the relative poses of several preceding and succeeding point cloud frames to generate local point clouds. Next, the Hidden Point Removal (HPR) operator [40] is applied to eliminate occlusion effects in the local point clouds. Finally, the local point clouds are projected into the event camera coordinate of the 8-mm and 16-mm lens camera pairs to obtain ground-truth depth.
@@ V SMALL-SCALE TESTBED
Hybrid Optical System Inivation DVXplorer DAHENG MER2 640$\times$480 1/3.5” FoV: 20°$\times$15° 1440$\times$1080 FoV: 17°$\times$13° 1/2.9” Rate: 25Hz
@@
Method CCRs1-low CCRs1-medium CCRs1-high CCRs2-low CCRs2-medium $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) STRTTC[32] 7.54 0.031 9.64 0.026 7.78 0.045 8.15 0.024 11.83 0.024 CMax[28] 2.56 3.833 3.44 3.701 7.39 3.228 5.02 2.32 11.20 3.449 ETTCM[30] 45.11 0.081 44.94 0.081 43.03 0.300 48.83 0.350 52.84 0.098 FAITH[27] 15.15 0.156 20.63 0.155 19.51 0.337 25.83 0.169 47.45 0.191 AEB-Tracker[33] 39.57 $\mathbf{4.2\times 10^{-5}}$ 42.41 $\mathbf{3.9\times 10^{-5}}$ 39.04 $\mathbf{1.1\times 10^{-5}}$ 37.09 $\mathbf{9.4\times 10^{-6}}$ 43.17 $\mathbf{7.3\times 10^{-6}}$ Image’s FoE[20] 5.37 0.036 5.12 0.031 1.86 0.041 3.76 0.031 3.85 0.031 Method CCRs2-high CCRm-low CCRm-medium Slider-750 Slider-1000 $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) $e_{\mathrm{TTC}}~{}(\%)$ Runtime (s) STRTTC[32] 13.68 0.028 9.87 0.031 10.36 0.023 8.95 0.015 12.43 0.016 CMax[28] 11.10 3.826 9.27 3.658 14.73 3.657 4.16 0.85 2.74 0.93 ETTCM[30] 49.46 0.127 71.05 0.059 58.00 0.121 18.99 0.498 15.20 0.191 FAITH[27] 49.62 0.165 75.78 0.116 56.26 0.11 34.45 0.186 47.94 0.233 AEB-Tracker[33] 42.43 $\mathbf{6\times 10^{-6}}$ 26.95 $\mathbf{1\times 10^{-5}}$ 35.99 $\mathbf{1.1\times 10^{-5}}$ 157.88 $\mathbf{1\times 10^{-5}}$ 165.56 $\mathbf{7.8\times 10^{-6}}$ Image’s FoE[20] 2.92 0.031 5.60 0.027 3.86 0.026 11.35 0.013 7.55 0.012